In a study of 258,001 cell phone​ users, it was found that 43 developed cancer of the brain or nervous system. Assuming that cell phones have no​ effect, there is a 0.000234 probability of a person developing cancer of the brain or nervous system. We therefore expect about 61 cases of such cancer in a group of 258001 people. Estimate the probability of 43 or fewer cases of such cancer in a group of 258,001 people.

What do these results suggest about media reports that cell phones cause cancer of the brain or nervous​ system? ​

(a) ​P(xless than or equals43​)equals nothing ​(Round to four decimal places as​ needed.)

​(b) What does the result from part​ (a) suggest about the media​ reports?

A. The media reports appear to be correct because one would expect that more than 43 cell phone users would develop cancer and the study offers significant evidence to support this.

B. The media reports appear to be correct because one would expect that less than 61 cell phone users would develop cancer and the study offers significant evidence to support this.

C. The media reports appear to be incorrect because one would expect that more than 61 cell phone users would develop cancer. In​ fact, the study may offer significant evidence to suggest that cell phone use decreases the probability of developing cancer.
Answer #1

a) P(x < 43)

= P(p < 43/258001)

= P(p < 0.000167)

= P(z < -2.236)

= 0.0127

b) Option C is correct.